(CO 3) Consider the table below Age Group Frequency 18-29 9831 30-39 7845 40-49 6869 50-59 6323 60-69 5410 70 and over 5279 What is the probability that the next person chosen will be in the 40-49 or 50-59 age groups?

42.5%

23.7%

35.4%

31.7%

Question 22 2 pts (CO 3) When children roll a die for a game, their probability of rolling a 5 is 17%. This is an example of which type of probability?

Classical probability

Public probability

Empirical probability

Subjective probability

Question 23 2 pts (CO 3) Consider the table below. Age Group Frequency 18-29 9831 30-39 7845 40-49 6869 50-59 6323 60-69 5410 70 and over 5279 What is P(over 60’)? 64.6%

71.8%

74.3%

25.7%
Answer #1

Answer

Question.CO3

total frequency = 9831+7845+6869+6323+5410+5279 =41557

the probability of being in age group 40-49 and 50-59 ={[ 6869/41557]}+{[6323/41557]}=0.1653(approx)+0.152152 (approx.)

=0.3174

0.3174×100= 31.74%

Question.No.22

Classical Probability

Question.No.23

Total frequency = 983 + 784 + 686 + 632 + 541 + 527 = 4153

Hence,

Probability of P(over 60)

= 5410/41557 +5279/41557

= 0.2572

25.72%